The toy factory produces X cast-iron toy aeroplanes per month at a fixed cost of R25 000. Each aeroplane costs R15 to produce and it is sold for R45 each. The total cost and profit functions are [1] T C = 15X + 25 000; π = 30X − 25 000. [2] T C = 45X + 25 000; π = 30X − 25 000. [3] T C = 45X; π = 60X − 25 000. [4] T C = 25 000 + 15X; π = 30X + 25 000. 


1
Expert's answer
2021-05-31T15:14:35-0400

Cost Function: Total cost of producing the units:

"C (X)= 25 000 + 15X"

Profit Function: Total Income minus Total Cost

"\pi(X)=45X-15X-25000=30X-25000"


Answer: [1] T C = 15X + 25 000; π = 30X − 25 000